What is the mass, in grams, of each of the following quantities of matter?

(a) 0.92 mole of copper atoms g
(b) 0.55 mole of nickel atoms g
(c) 2.03 moles of nitrogen atoms g
(d) 2.24 moles of nitrogen molecules (N2) g

How many atoms are present in the following molar quantities of various elements?

(a) 1.85 moles Fe atoms Fe
(b) 1.95 moles Ni atoms Ni
(c) 1.45 moles C atoms C
(d) 1.45 moles Ne atoms Ne
Answer #1

a) 0.92 mole of copper = mole * molar mass = 0.92 mole * 63.5 g / mole = 58.42 g

b) 0.55 mole of nickel atoms = 0.55 mole * 58.69 g / mole = 32.28 g

c)  2.03 moles of nitrogen atoms = 2.03 mole * 14 g / mole = 28.42 g

d) 2.24 moles of nitrogen molecules (N2) = 2.24 mole * 28.0 g / mole = 62.72 g

1 mole atom = 6.023 * 10^23 atoms

a) 1.85 moles Fe = 1.85 * 6.023 * 10^23 = 1.11 * 10^24 Fe atoms.

b)  1.95 moles Ni = 1.95 * 6.023 * 10^23 = 1.17 * 10^24 Ni atoms.

c)  1.45 moles C = 1.45 * 6.023 * 10^23 = 8.73 * 10^23 C atoms.

d) 1.45 moles Ne = 1.45 * 6.023 * 10^23 = 8.73 * 10^23 Ne atoms.